### v1

# Artificial Neural Networks for Magnetoencephalography: A review of an emerging field

Abstract
Magnetoencephalography (MEG) is a cutting-edge neuroimaging technique that measures the intricate brain dynamics underlying cognitive processes with an unparalleled combination of high temporal and spatial precision. MEG data analytics has always relied on advanced signal processing and mathematical and statistical tools for various tasks ranging from data cleaning to probing the signals’ rich dynamics and estimating the neural sources underlying the surface-level recordings. Like in most domains, the surge in Artificial Intelligence (AI) has led to the increased use of Machine Learning (ML) methods for MEG data classification. More recently, an emerging trend in this field is using Artificial Neural Networks (ANNs) to address many MEG-related tasks. This review provides a comprehensive overview of how ANNs are being used with MEG data from three vantage points: First, we review work that employs ANNs for MEG signal classification, i.e., for brain decoding. Second, we report on work that has used ANNs as putative models of information processing in the human brain. Finally, we examine studies that use ANNs as techniques to tackle methodological questions in MEG, including artifact correction and source estimation. Furthermore, we assess the current strengths and limitations of using ANNs with MEG and discuss future challenges and opportunities in this field. Finally, by establishing a detailed portrait of the field and providing practical recommendations for the future, this review seeks to provide a helpful reference for both seasoned MEG researchers and newcomers to the field who are interested in using ANNs to enhance the exploration of the complex dynamics of the human brain with MEG.
Keywords Artificial Neural Networks (ANNs)- Magnetoencephalography (MEG) - Deep Learning (DL) - Brain Imaging, Machine Learning (ML) - Brain Decoding
1 Introduction
Artificial Neural Networks (ANNs) have had an enormous impact on most research fields, especially in computer vision and natural language processing ([1]). In recent years, ANNs have become increasingly used for brain data analyses and modeling, yielding ample evidence for their added value in neuroscience research ([2, 3]). Deep Learning (DL) has been applied in the analysis of various types of brain imaging modalities, including structural Magnetic Resonance Imaging (MRI), functional Magnetic Resonance Imaging (fMRI), Electroencephalography (EEG), and Magnetoencephalography (MEG). While using ANNs specifically with EEG has been the subject of a comprehensive review (e.g. [4]), using ANNs with MEG data is less common, and the current state-of-the-art has yet to be surveyed. The present review seeks to fill in this gap by providing a detailed survey of the main applications of ANNs in MEG research to date. The strengths and limitations of the approaches reviewed allow us to highlight the potential of ANNs in future MEG research and the main challenges in this field.
1.1 Magnetoencephalography and brain imaging
MEG is a non-invasive electrophysiological brain imaging technique that detects the magnetic signals generated by the electrical activity generated by synchronized neuronal ensembles in the brain. Although it shares essential similarities with EEG (such as the electrophysiological origins of the signals and the millisecond-range temporal resolution), MEG offers unique advantages. A significant merit of MEG over EEG lies in its lesser susceptibility to spatial smearing caused by variations in the conductivity of different brain tissues and the skull. MEG is sensitive to magnetic permittivity changes, which are relatively homogeneous in the brain ([5]). As a result, MEG is a better-suited technique for source reconstruction and the analysis of spatio-temporal brain dynamics in cortical source space. Compared to fMRI, another widely used functional neuroimaging technique, MEG has several strengths. While the latter generates images of the brain only every few seconds, MEG captures brain activity on a millisecond scale. Furthermore, unlike fMRI, which measures variations in oxygenation levels in the brain, MEG provides a direct measure of the brain’s electrophysiological activity [6]. Both EEG and MEG produce sensor-level data that are highly intercorrelated and have limited spatial interpretability. These ambiguities can be reduced with MEG by applying source localization techniques ([7, 8]). Although source reconstruction does not entirely resolve the field spread problem or linear mixing issue, it does facilitate the exploration of the functional role of specific brain areas with more anatomical precision than what can be achieved with sensor-level analysis [9]. Overall, given its high temporal and spatial resolution, MEG has become an established imaging modality considered particularly suited for addressing neuroscientific inquiries involving intricate spatial, temporal, and spectral patterns of distributed brain dynamics. These include, for instance, the real-time integration of information in the brain from one or multiple modalities and the neural dynamics that underpin advanced cognitive functions such as language processing and decision-making. For a complete overview of the advantages and drawbacks of MEG in the field of neuroscience, see [10].
1.2 Standard M/EEG analysis pipeline
Generally, the M/EEG analysis pipeline consists of a sequence of processing steps that start from the sensor-level raw data and, after a series of manipulations, ultimately yield observations that address a research question. Standard preprocessing pipelines often include data cleaning and artifact rejection procedures, including channel rejection, notch-filtering, and Independent Component Analysis (ICA). M/EEG recordings yield rich and high-dimensional data, and typically, the analysis pipeline involves computing a set of variables from the raw data, also known as hand-crafted feature extraction in the case of ML analysis. Computing these variables (or features) relies on processing pipelines that can be broadly divided into two main categories: time-domain or frequency-domain analyses (which can also be jointly explored using time-frequency representations of the data). Time domain analyses typically involve the computation of time-based features such as event-related potentials (ERPs), long-range temporal correlations (LRTC) using Detrended Fluctuation Analysis (DFA), complexity measures (e.g. Hjorth parameters, Hurst exponent), fractality measures (fractal dimension) or entropy measures (approximate entropy, sample entropy, permutation entropy). Frequency domain analyses derive measures from the frequency domain using spectral analysis (e.g. Wavelet and/or Fourier transform), with typical features including the spectral power across canonical frequency bands (e.g. delta, theta, alpha, beta, gamma). In time- or frequency-domain analyses, connectivity features can be calculated using multi-channel or whole-brain source data. Typical connectivity metrics assess different associations between signals from various sources or sensors. These metrics include magnitude-squared coherence (MSCoh), Phase Locking Value (PLV), Phase-Lag Index (PLI), and weighted PLI (wPLI), to name but a few. Different features capture complementary -although sometimes partly overlapping- aspects of the neural dynamics associated with various behavioral states. Sensor-level analyses in neuroscience are informative, but source reconstruction is essential for improving the interpretation of the spatial distribution of significant effects. This step involves defining a forward model and solving the inverse problem. Effective source reconstruction allows for precise identification of neural markers in specific brain regions, enhancing our understanding of complex brain functions and informing both hypothesis validation and exploratory data-driven research.
1.3 Machine learning versus inferential statistics
A typical M/EEG neuroscience study employs either inferential statistics, using hypothesis testing methods such as t-tests or ANOVA, or statistical learning approaches, notably machine learning (ML). The choice between these approaches depends on the specific research questions, with inferential statistics often used to test predefined hypotheses about neural activity, while ML is typically applied to explore complex patterns and predictions from high-dimensional datasets. Both approaches have significantly contributed to progress in neuroscience, as detailed in [11]. Additionally, the Bayesian framework has emerged as a complementary avenue in M/EEG analysis [12], allowing hypothesis testing from a probabilistic perspective. While hypothesis testing, whether conducted through traditional methods or Bayesian approaches, has long been a staple in neuroscience, it remains the subject of ongoing debate and scrutiny (e.g. [13]). Despite its undeniable strength and utility, one significant critique of null hypothesis significance testing (NHST) is the fact that it strongly relies on a priori choices guided by the literature and expert domain knowledge and that generalization of the findings is merely assumed on the basis that the studied sample is representative of the target population. As an alternative approach to examining hand-picked variables informed by the literature, MEG researchers have begun to use a variety of data-driven approaches. These broadly include two types of approaches to relevant feature identification with hardly any a priori information: (a) Feeding a large number of hand-crafted features into a decoding framework and assessing their relative contributions to model performance, and (b) Learning the relevant features from the data themselves, a form of data-driven automatic feature extraction known as representation learning. Unlike simpler algorithms such as Support Vector Machines (SVM), Linear Discriminant Analysis (LDA), or Decision Trees and Forests, ANNs can learn and extract useful features from the data automatically through intricate computations, resulting in an optimal data representation tailored to the specific requirements of the task. Despite its conceptual appeal, representation learning comes with its limitations, in particular when it comes to interpreting the learned features. ANNs are often criticized for being ’black boxes’ compared to more classical ML algorithms, which use hand-crafted features as inputs. Enhancing the feature interpretability is an active research topic in deep representation learning.
1.4 Artificial Neural Networks in a nutshell
Understanding Artificial Neural Networks (ANNs) begins with grasping the concept of linear regression, where the output $y$ is given by $y=wx+b$, with $w$ representing the weight, $x$ the input, and $b$ the bias. In ANNs, this equation gets extended by an activation function $f$, transforming it to $y=f(wx+b)$. As the neural network gains depth through additional layers, these equations combine to form increasingly complex mathematical models. A neural network’s complexity, or capacity, indicates its ability to model intricate functions, but this can be a double-edged sword. High capacity often leads to overfitting, where the model learns the training data too well, capturing noise and inaccuracies, thus performing poorly on new data. To mitigate overfitting, regularization techniques such as Maxpooling, which reduces feature map dimensions, Dropout, which nullifies a subset of neurons during training, and Batch Normalization, which standardizes layer outputs, can be applied. Another regularization method comes in the form of Data Augmentation, where the initial dataset can be enlarged by introducing noise to the data, rotating, or cropping it. Learning in ANNs occurs through optimization, typically employing Stochastic Gradient Descent (SGD). During each training iteration, the model’s parameters are adjusted based on computed gradients from a loss function, facilitated by a technique known as backpropagation. As we progress to more advanced ANNs, we encounter Convolutional Neural Networks optimized for image data, Recurrent Neural Networks ideal for sequences, and Attention Models that allocate varying degrees of focus on different parts of the input. Overall, the domain of Artificial Neural Networks (ANNs) extends from fundamental mathematical principles to intricate architectures and sophisticated methods for effective training.
1.5 ANNs for neuroimaging data
Recently, ML has garnered much interest from the neuroscience research community ([14, 15, 16]). While deep learning has predominantly been applied to fMRI and MRI data due to their compatibility with existing architectures, its use in EEG analysis is also growing, partly due to EEG’s wider availability. However, the unique strengths of MEG, such as its high temporal and spatial resolution, present a compelling case for its integration with advanced techniques like ANNs. These networks are well-suited to handle MEG’s complex, high-dimensional data. They apply to various tasks, including classification, data modeling, representation learning, data cleaning, and source estimation, thereby contributing to a more nuanced understanding of brain functions. The use of ANNs to analyze MEG data is gaining traction, offering novel perspectives and data analyses that standard MEG methodologies typically do not provide. Although arguably still in the early stages, exciting progress has already been reported in this emerging field.
1.6 Goal of this review
This review provides a comprehensive survey of ANN applications in magnetoencephalography (MEG) data analysis, highlighting their benefits and potential added value. These techniques have proven to be remarkably effective in various neuroscience contexts. By providing examples of functional architectures and techniques, we aim to provide an up-to-date overview of the concrete progress in ANN-based MEG studies. We hope to expedite researchers’ path to developing their own implementations. More globally, we hope this review will encourage researchers well-versed with MEG to integrate deep learning techniques into their work, and conversely encourage experts in deep learning with an interest in neuroscience to consider MEG as a particularly promising brain measurement modality in the context of Neuro-AI research applications.
2 Methods
2.1 Litterature research
For this review, we collected English-language conference and journal papers using PubMed, Google Scholar, arXiv, and bioRxiv. We used the following query to find the papers referenced in this review: ("deep learning" OR "artificial neural network*" OR "Convolutional neural net*" OR CNN OR "Recurrent neural net*" OR RNN) AND ("MEG" OR Magnetoencephalogra*) Relevant papers were selected by examining the title first, followed by the abstract. Finally, we used a PDF search for the terms from the query to better understand how they are used. This study includes conference papers, journal papers, and pre-prints. The final inclusion criterion requires papers to possess a valid Digital Object Identifier (DOI) and a publication date preceding November 2024.
Tables 3, 4, 5, 6 provide an overview of the key information extracted from the included studies for this survey. More detailed information on each paper was collected in a table we have made available online as a Google sheet: https://tinyurl.com/ub3s5mr. See Table 1 for details on each information category extracted from the papers.
2.2 Types of studies
In conducting this review, we opted to categorize the included studies into three distinct groups: (1) Classification, (2) Modeling, and (3) Other. The ’Classification’ category was split into four sub-groups: Decoding, Brain-Computer Interfaces (BCI), Clinical Applications or Computer-Aided Diagnosis (CAD), and Event Detection. Studies under ’Decoding’ primarily leverage classification as a statistical tool to identify the involvement of specific brain features in certain tasks. The BCI subcategory includes studies that incorporate ANNs in their BCI framework. Meanwhile, the ’Clinical’ subcategory includes studies that utilize ANNs to enhance clinical diagnosis or prognosis, and the ’Event Detection’ subcategory regroups studies using ANNs to continuously detect events in MEG data. Studies in the ’Modeling’ category primarily focus on comparing activations across ANNs and the brain. The ’Other’ category gathers the studies that do not fit the aforementioned groups and includes research utilizing ANNs for preprocessing tasks such as artifact detection and removal or source localization (See Table 2 for further details).
Item Description Online table Included table
Author Name of the main author(s).
Publication year The year the paper/pre-print was published.
Title The Title of the paper.
DOI The DOI of the paper.
Category The paper category (classification, modeling, or other).
Subcategory Subcategory of paper.
Goal The goal of the study.
Subjects (N) The number of subjects for the study.
Notes on subjects Information and details on subjects
Sample size (trials) The total number of samples used.
Notes on sample size Information and details on the number of samples.
Input data The type of input data given to the neural network.
Detailed input data A more detailed version of the input data column.
Data augmentation Data augmentation techniques used, if any.
Validation The type of validation or model selection technique used.
Sensors The number of sensors used.
SF The sampling frequency used.
Performance The performances that the paper was able to reach.
Trial length Length of a single data point given to the Neural Network.
Baseline Previously attained/baseline performances.
Architecture (depth) The type of architecture used and its depth.
Detailed architecture The detailed architecture of the network used.
preprocessing The detailed preprocessing steps used.
Processing Data processing/feature extraction process.
Specificities Further details on how the study was led.
Interpretation visualization techniques or interpretation tools, if any.
Training params Details on the neural network training, when available.
Code availability Whether the code is publicly available or not.
Table 1: Description of table items, and in which table they can be found. Full table and information can be found in https://tinyurl.com/ub3s5mr
2.3 Notes on terminology
Some of the ML nomenclature overlaps with terminology used in M/EEG research, which can sometimes lead to confusion. In particular, the term ’epoch’ in deep learning generally describes a complete pass of the training data. Meanwhile, in the field of M/EEG, it usually describes a segment of data (or a single trial). In addition, the term ’samples’ in ML usually refers to individual examples from the dataset, which should not be confused with the use of the same term to refer to a data point in a M/EEG time series, where the parameter known as ’sampling rate’ (Hz) refers to the rate at which the data is assessed, i.e., data points per second.
3 Results
In the following, we present a comprehensive analysis of the corpus of papers selected for this review. After a general overview highlighting key trends and patterns across various categories, we dive into a more detailed examination of the methods, data, and ANN architectures used across each of the three main categories of papers.
The methodology explained in the previous section (2) led to the identification of 119 relevant studies. Among the papers considered, 102 were peer-reviewed, while 17 had not yet undergone peer review at the time of writing. Our categorization method resulted in the distribution of 69 papers under ’Classification,’ accounting for 58% of the total; 16 papers were categorized as ’Modeling,’ making up 13.4%; and 34 papers fell into the ’Other’ category, representing 28.6% of the study corpus included in this review (See Fig 1). The category containing most of the papers is the ’Classification’ category, followed by ’Other’; the category regrouping the least amount of studies is ’Modeling’. Generally, the trend of using ANN on MEG data has increased over the years. A similar growth trajectory is anticipated for studies on this topic, analogous to the rapid expansion of research utilizing Artificial Neural Networks (ANNs) for Electroencephalography (EEG) data analysis ([17]).
3.1 General Overview
The in-depth survey revealed several key observations about this emerging field (Fig 1). Interestingly, out of the 119 selected studies for this review, 64 use less than 20 subjects, 34 use between 20 and 100 subjects, and 15 use more than 100 subjects. The remaining studies use synthetic MEG data or do not mention the number of subjects, or patients included in the study. The number of samples (number of trials or epochs) in datasets varies across a wide range, from 17 to 150000, with an average of 76261, a standard deviation of 289279, and a median of 3540. However, a few outliers significantly impact the average and standard deviation of the number of samples used in the included studies (Correcting for outliers gives an average sample number of 13249 and a standard deviation of 21247). Furthermore, the length of the data segments fed to the neural network varies between 300ms to 10s. The most commonly used sampling frequencies are 1000Hz (26 papers) and 250Hz (16 papers), and the median across all studies was 600Hz.
[…]
style=plain,subcapbesideposition=top
[…]
[(a)]Refer to caption
[…]
[(b)]Refer to caption
[…]
[(c)]Refer to caption
[…]
[(d)]Refer to caption
Figure 1: Quantitative overview of the reviewed studies (n=119). (a) The evolution of the number of MEG publications using ANNs across categories (classification, modeling, and other). (b) Size of the dataset (number of samples) used for training and evaluation of the ANNs across the reviewed studies. ’N/M’ accounts for the studies that did not mention the number of samples used. (c) Overall distribution across categories of all the reviewed publications. (d) Overall distribution across subcategories categories of all the reviewed publications.
Category Subcategory Study
Classification Decoding [18], [19], [20], [21], [22], [23], [24], [25], [26], [27],
[28], [29], [30], [31], [32], [33], [34], [35], [36],[37],
[38], [39], [40], [41], [42], [43], [44], [45]
BCI [46], [47], [48], [49], [50], [51], [52], [53],
[54], [55], [56], [57], [58], [59]
Clinical [60], [61], [62], [63], [64], [65], [66], [67], [68], [69],
[70], [71], [72], [73], [74], [75], [76]
Event detection [77], [78], [79], [80], [81], [82], [83], [84], [85][86]
Modeling Visual Cortex [87], [88], [89], [90], [91], [92], [93], [94], [95] [96], [97]
Auditory Cortex [98], [99], [100], [101], [102], [103]
Other Preprocessing [104], [105], [106], [107], [108], [109], [110], [111],
Methods [112], [113], [114], [115], [116], [117], [118], [119],
[120], [121], [122], [123], [124], [125], [126], [127]
Source Localization [128], [129], [130], [131], [132], [133], [134],
[135], [125]
Table 2: The categories and subcategories classification for all reviewed papers.
Year Authors Subcategory N trials Architecture (depth) Validation Code
2016 Yu et al Decoding 15 2000 CNN (3) train/test split
2017 Wang et al Decoding 2 410 MLP (1)
4Fold on train +
holdout test set
2018 Hramov et al Decoding 5 150000 MLP (1) train/valid split
2018
Frolov
and Pisarchik
Decoding 5 10000 MLP (1) train/test split
2019 Dash et al Decoding 4 N/M
MLP (3)
LSTM
train/valid/test split
2019 Garry et al Decoding 605 75396 CNN (3) train/valid/test split
2019 Kim et al Decoding 27 5400 CNN + Attention (6) 5Fold
2019 Kostas et al Decoding 101 3900 CNN and RCNN (N/A) 5Fold
2019 Dash et al Decoding 8 2400 MLP (1) train/test split
2019 Huang and Yu Decoding 16 588 CNN (4) train/test split
2020 Abdellaoui et al Decoding 18 N/M CNN + Attention + LSTM (10) train/test split
2020 Dash et al Decoding 7 300 MLP (1) 5Fold
2020 Dash et al Decoding 10 180 MLP (1) 5Fold
2021 Li et al Decoding 16 9414 2 x GRU + FC LOO
2021 Feng et al Decoding 200 397 GoogLeNet - CNN (144) train/valid split
2021 Chang et al Decoding 8 108 GAN-like (5) Holdout set
2021 Pilyugina et al Decoding 17 17 CNN (1) train/test split
2022 Engemann et al Decoding 646 N/M ShallowFBCSPNet + Deep4Net 5Fold
2022 Csaky et al Decoding 15 53100 WaveNet* train/valid split
2022 Shi et al Decoding 17 10880 EEGNet (4) nested LOSO
2023 Zhang et al Decoding 20 N/M N/M Leave One Out
2023 Csaky et al Decoding 15 53100 Wavenet Classifier (9) LOSO
2023 Özer et al Decoding 18 9414 LSTM, GRU and CNN 10Fold
2023 Bu et al Decoding 12 8640 CNN (4) 5Fold + train/test split
2024 Boyko et al Decoding 96 2304000 N/M train/valid/test split
2024 Yang et al Decoding 27 179977 N/A train/valid/test split
2024 Zubarev et al Decoding 12 3600 LF-CNN (3) 9Fold
2024 Yang et al. Decoding 27 179977 NeuGPT train/valid/test
Table 3: Part 1 of 4 of the table containing condensed information about the included papers. This part includes all papers categorized as ’Classification’ papers and subcategorized as ’Decoding’ papers. More info can be found in the complete table: https://tinyurl.com/ub3s5mr. N is the number of subjects included in the study. ’N/M’ is used when specific information was either not mentioned or when it was impossible to infer from other information. N/A Means the information does not apply to the study.
Year Authors Subcategory N trials Architecture (depth) Validation Code
2018 Meng et al Clinical 45 N/M CNN (4) 4Fold
2019 aoe et al Clinical 233 70000 CNN (12) 10Fold
2020 Gu et al Clinical 32 N/M MLP (4) + Self-attention head cross-validation
2020 Zhang et al Clinical 32 N/M LSTM (2) 10Fold + holdout test set
2021 Xu et al Clinical 129 N/M G2G network (1) train/valid split
2021 Giovannetti et al Clinical 87 6525 AlexNet - CNN (5) LOSO
2021 Wu et al Clinical 38 2280 CNN (4) 2Fold
2021 Huang et al Clinical 95 95 EEGNet (4) train/valid/test split
2022 Huang et al Clinical 190 33500 ResNet (5) 5Fold + train/val/test split
2022 Fujita et al Clinical 180 21000
MLP (1)
MNet (12)
10Fold + holdout test set
2023 Barik et al Clinical 60 N/M MLP (1) 5Fold nested
2024 Anand et al Clinical N/M 1080000 N/M train/test split
2024 Achterberg et al Clinical 8 2400 DNN (4) train/valid/test split
2018 Dash et al BCI 4 1635 MLP (1) train/valid/test split
2018 Dash et al BCI 3 1225 CNN (8) 5Fold (trials)
2019 Hramov et al BCI 10 N/M MLP N/M
2019 Hramov et al BCI 7 N/M MLP (2) N/M
2019 Zubarev et al BCI 7 11354 LF-CNN (3) LOO
2019 Dash et al BCI 4 1635 CNN (5) train/valid/test split
2020 Dash et al BCI 8 3046
AlexNet (7)
ResNet101 (101)
Inception-ResNet-v2 (164)
train/valid/test split
2020 Dash et al BCI 8 3046 LSTM (3) N/M
2020 Yeom et al BCI 9 2160 LSTM 5Fold
2020 Dash et al BCI 4 1500 LSTM* train/valid/test split
2020 Lopopolo et van den Bosch BCI 15 1377 CNN (4) houldout test set
2021 Ovchinnikova et al BCI 32 422 LF-CNN (3) nested 4Fold in a 5Fold
2022 Fan et al BCI 61 39040 CNN (3) train/test split
2018 Guo et al Event 10 102 SSAE (3) 5Fold
2019 Zheng et al Event 20 4000 CNN (8) multiple KFold
2020 Liu et al Event 20 150 CNN* KFold
2021 Zhang Chuncheng et al Event 10 N/M EEGNet (4) cross-validation
2022 Hira et al Event 348 23177 ResNet (26) and AE 5Fold and 10Fold
2022 Bhat et al Event 15 11000 N/M 5Fold
2022 Zhao et al Event 20 1320 DANN Leave One Out
2022 Guo et al Event 20 202 Transformer KFold
2022 Zhang et al Event 20 150
CADNet
DendriteNet
N/M
2023 Zheng et al Event 48 N/M CNN (5) 7Fold
2023 Mouches et al Event 95 40662 N/M 10Fold
2024 Wei et al Event 277 230325 N/M train/valid split
2024 He et al Event 11 35013 CNN + Attention (5) LOSO
2024 Dev et al Event 7 7060 N/M train/valid/test split
2024 Hira et al. Event 1782 60139 26 Layer scSE-ResNet 5Fold subject-wise
Table 4: Part 2 of 4. Includes all papers categorized as ’Classification’ papers and subcategorized as ’Clinical’, ’BCI’, or ’Event Detection’ papers.
Year Authors Subcategory N trials Architecture (depth) Validation Code
2016 Cichy et al Visual 15 3540 SuperVision - CNN () N/M
2017 Cichy et al Visual 15 N/M CNN (8)
Bootsrapping
resampling
2017 Seeliger et al Visual 15 1000 VGG-S - CNN (10) 10Fold
2018 Dima et al Visual 19 720
CNN (7)
AlexNet - CNN (8)
5Fold
2018 Bankson et al Visual 32 32 VGG-F - CNN (7) LOSO
2019 Rajaei et al Visual 15 1536
AlexNet - CNN (7)
HRRN (152)
LOO
2019 Kietzmann et al Visual 15 380
CNN
RCNN (6)
2Fold
bootstrapping
2020 Giari et al Visual 25 3525 AlexNet - CNN (5) N/A
2022 van Vliet et al Words 15 560 VGG (11) N/A
2023 von Seth et al Visual 36 N/M N/M N/M
2019 Donhauser Speech 11 77 LSTM* 7Fold
2022 Caucheteux and King Auditory 92 N/M
Transformer
CNN (4, 8 or 12)
N/A
2022 Wingfield et al Speech 16 66300 MLP (6) N/A
2023 Desbordes et al Speech 11 3630
LSTM (2)
Transformer (11)
10Fold
2024 Brodbeck et al
Speech
Recognition
18 N/M N/M N/M
2024 Lyu et al
Speech
Recognition
16 5760 N/M N/M
Table 5: Part 3 of 4. Includes all papers that were categorized as ’Modeling’ papers.
Year Authors Subcategory N Trials Architecture (Depth) Validation Code
2019 Dinh et al SL 1 N/M LSTM train/valid split
2021
Pantazis
and Adler
SL N/A N/M CNN (5) and MLP (4) N/A
2021 Dinh et al SL 1 1653 LSTM train/valid split
2022 Sun et al SL 29 620256 ResNet (4) LSTM (3) train/test split
2023 Sun et al SL 29 620256 N/M N/M
2023 O’Reilly et al SL 1 446 SimpleRNN (4) Bootstrapping
2024
Sanchez-Bort
et al
SL 1 N/M MPSS (4) N/M
2024 Jiao et al SL 1 N/M CNN + Attention N/M
2024 Yokoyama et al. SL 3 N/M 4LCNN (4) train/valid sets
2016
Hyvärinen
and Morioka
Preprocessing 9 N/M FC (4) M/M
2017 Garg et al Preprocessing 49 980 CNN (6) train/test split
2017 Garg et al Preprocessing 44 880 CNN (9) LOO + train/test split
2018 Croce et al Preprocessing 67 4038
CNN (6); FC (6)
CNN+FC
10Fold (trials)
2018 Hasasneh Preprocessing 48 1112
temporal CNN (3)
spatial CNN (2)
sCNN + tCNN
50Fold
2021 Feng et al Preprocessing 4 66780 GoogLeNet - CNN (144) train/valid/test split
2021 Treacher et al Preprocessing 217 294 CNN (11) 10Fold
2023 Hamdan et al Preprocessing N/M N/M AE (9) N/M
2017 Guo et al Method 3 N/M CNN (5) 10Fold
2019 Harper et al Method 72 N/M LRCN train/valid/test split
2020 Shu and Fyshe Method 9 60 SAE LOO
2021 Abdellaoui et al Method 18 N/M CNN + Attention train/test split
2022 Priya and Jayalakshmy Method 23 13472 GoogLeNet
KFold with
holdout test set
2023 Gosti et al Method 10 N/M RHoMM - RNN (N/M) N/M
2023 Elshafei et al Method 1 166800
DNN (4)
CNN (1)
train/test split
2023 Fan et al Method 669 70000
LFCNN
VARCNN
HGRN
train/valid split
2023 Csaky et al Method 37 53100 WaveNet-based train/valid split
2023 Zhu et al Method 676 100000
MLP (1)
NICA(TCL) (3)
NICA(IIA) (3)
N/A
2024 Solana et al Method 16 N/M
ROCKET-based
models (3)
train/valid split
2024 Fan et al Method 250 29238
EEGNet
ResNet
ShuffleNet
N/M
2024 Gallard et al Method 44 N/M CycleGAN N/A
2024 Chou et al Method 17 N/M CNN (3) train/valid/test split
2024 Csaky et al Method 15 3540 GPT train/valid/test split
2024 Gideoni et al. Method 6 N/M MLP and CNN train/valid/test split
2024 Ferrante et al. Method 4 N/M CNN and CLIP N/A
Table 6: Part 4 of 4. Includes all papers that were classified as ’Other’ papers. SL is an abbreviation of Source Localization.
3.2 Classification studies
As explained in section 2, we split the studies in this category into three MEG-based classification problems: Decoding problems, Clinical studies, and Brain-Computer Interfaces (BCI).
Decoding studies: Generally speaking, decoding studies aim better to understand the neural correlates of behavior and cognitive processes. This is typically accomplished by using ML or statistical inference. This approach determines the neural responses that underlie specific sensory, motor, or higher-order cognitive functions by revealing the features that exhibit the strongest separation between classes. In this category, 14 studies across four groups of researchers use ANNs for speech decoding or speech-related BCI ([29, 46, 51, 50, 23, 21, 48, 49, 22, 24, 18, 58, 38]). Three studies specifically focus on using ANNs for auditory stimuli decoding ([36, 20, 34]). Additionally, 10 studies use them for visual stimuli decoding ([25, 26, 27, 28, 31, 33, 34, 42, 44, 40]). The remaining eight decoding studies use ANNs for affect states decoding ([30]), age decoding ([43]), hand gestures ([37, 35]), rhythm decoding ([19]), text decoding ([41, 45]), or decoding methods ([32]).
Clinical studies: The general goal of studies classified as Clinical is to predict or detect diseases or anomalies in patients. Thirteen studies use ANNs for clinical purposes or computer-aided diagnosis (CAD). Two focus on epilepsy-related signal properties detection ([71, 66]). Two studies focus on Alzheimer’s disease early detection [62] or evolution [68]. Three focus on neurological disease diagnostic ([60, 70, 69]), and the remaining studies focus on schizophrenia detection [67], autism detection in children population [72], PTSD severity evaluation [65], migraine diagnostic [61], Mild traumatic brain injury detection [63] or depression and bipolar disorder detection [64]. Although not classified as clinical studies, among the 15 studies sub-classified as ’Event Detection’, 13 are focused on detecting and/or locating epileptic spikes ([78, 77, 84, 85, 82, 79, 80, 83, 74, 76, 73, 75, 86]). Therefore a combined total of 14 studies out of the 69 classification studies are epilepsy-related.
BCI studies: The third subcategory regroups the BCI studies, which generally assess the feasibility of building brain-controlled devices that rely on predicting the subject’s intentions from their neural data. BCI studies cover a wide range of approaches. Still, most consist either of offline assessments of previously recorded data or online examination of intention decoding in a closed-loop setting. BCI is one of the neuroscience domains in which ML made its earliest incursions (e.g. decoding arm movement direction and kinematics from neuronal activity in non-human primates). Of the remaining six studies subcategorized as BCI, five use ANNs for motor tasks ([52, 53, 56, 55, 54]), and finally, in [57], the authors investigate how ANNs can improve the state-of-the-art of BCI.
Standard ML algorithms (e.g. support-vector machine, linear discriminant analysis, or random forest) are increasingly used in MEG-based decoding studies. However, according to our survey, the first ANN-based MEG classification studies were published in a journal only in 2016 [30]. To date, the MEG studies that have applied ANNs for classification purposes are numerous and very different from one another. In particular, due to differences in the goal and rationale, the 69 classification studies vary substantially in terms of number of subjects (ranging from 2 [29] to 646 [43]). However, most studies, 21 of the 56 that mentioned it, use more than 1000 samples in their dataset, as seen in Fig. 1. Thirteen studies in this category did not mention the number of samples. Seventeen papers in this category have their codes publicly available or available upon request. Ten of these studies were published in 2023 or later. We will now provide an overview of the datasets used, followed by an overview of the network architectures implemented by the various papers.
3.2.1 The data
The studies in this group used MEG data sampled at frequencies between 50Hz and 2400Hz. Still, nine studies omitted details about the sampling frequency, in most cases because the authors extracted features directly from the raw signals, making the original sampling rates less pertinent([61, 62, 72, 71, 30, 114, 115, 53, 19]).
Typically, across all studies in this category, the MEG data were preprocessed before being fed as input to a neural network. This preprocessing phase generally includes down-sampling, band-pass filtering, and either de-noising and/or artifact removal. Of the 69 studies in this category, 12 studies (Decoding: [21, 22, 23, 29] BCI: [46, 47, 48, 49, 51, 53], Clinical: [64], Event Detection: [85]) use Discrete Wavelet Transform to increase Signal-to-Noise Ratio (SNR) in the data as a preprocessing step. Notably, these 12 studies are spread across only four distinct research groups representing a small portion of the compendium. Five studies employ alternative de-noising techniques ([28, 42, 69, 70, 77]). Additionally, 12 studies use ICA to remove heart and/or eye movement artefacts ([34, 37, 44, 57, 72, 67, 78, 81, 69, 70, 35, 40]). Three studies use Fieldtrip artefact removal algorithm ([68, 61, 62]) or EEGlab eye artefact removal ([18]). The remaining 38 studies do not explicitly mention de-noising or artifact removal techniques in their preprocessing pipeline.
Among all classification papers, 27 studies make use of hand-crafted features. While most of these studies train the network uniquely with hand-crafted features ([53, 61, 46, 47, 66, 51, 62, 63, 68, 75, 73, 73, 71, 23, 30, 22, 24, 29, 33, 64, 70, 59, 85]), a few use a combination of hand-crafted features and raw MEG data ([65, 20, 60, 69]. Among the reviewed studies, spectral power and functional connectivity matrices were the two most frequently used types of hand-crafted features.
Data augmentation is often necessary when using ANNs, especially considering that small datasets can hinder generalization performances and that acquiring MEG data is expensive. Among the surveyed classification studies, 12 make use of augmentation techniques ([76, 46, 23, 18, 64, 39, 42, 40, 70, 38, 45, 86]). These studies use some form of linear shifting and/or sensor shuffling or mix-up regularization ([136]). An example of linear shifting consists of selecting a different cropping of a trial or segment around the cue in the temporal dimension. The shifted samples end up having the same length but different onsets.
In the studies that used the original MEG data as input rather than hand-crafted features, the reported length of the segments used as samples for training and testing the ANNs varied between 140 and 2,400. Unfortunately, a few papers do not mention the length of the segments used as input for their neural network ([26, 31, 19, 69]).
3.2.2 The networks
The most commonly used ANNs for classification-type studies are Convolutional Neural Networks (CNNs). The second most used architecture was Multi-Layer Perceptrons (MLP), followed by Recurrent Neural Network (RNN) architectures. For the CNNs, the depth of the networks varied between one and 144 hidden layers. The papers that used MLPs essentially limit themselves to six hidden layers.
The CNNs were designed to mimic the human visual cortex. It comprises successive convolution layers that can learn abstract representations of the data. The convolution layers are generally followed by two to three Fully Connected (FC) layers that will perform the final classification. Long Short-Term Memory Networks (LSTMs) are an advanced form of Recurrent Neural Networks (RNNs), addressing some issues with long-term dependencies in the original RNN design. RNNs can learn sequential/temporal relationships in data and predict the next element of a sequence. Autoencoders are MLP architectures where the first half of the networks consists of Fully Connected (FC) layers of decreasing sizes, followed by FC layers of increasing sizes up to the size of the input. Such architecture aims to learn an ideally lossless compressed representation of the data. Sparse Autoencoders (SAEs) are a special case of AEs where a sparsity constraint is added to the hidden units: most units composing the hidden FC layers will randomly be set to zero for any given input. More info on these networks can be found in [1].
Like any other parametric algorithm, ANNs have parameters that significantly influence their ability to learn and acquire a good data representation. These parameters define the architecture’s depth and shape, and the optimal choices vary depending on the application. This section provides an overview of the predominant architectures and parameter configurations employed in the studies categorized under ’classification’.
Most of the CNN architectures used in the reported MEG classification papers comprise the input layer followed by multiple blocks convolution layers with maxpool and/or Rectified Linear Unit (ReLU) layers. At the end, one or two fully connected layers are followed by an activation function. For the convolution layers, the critical parameters are filter size, stride, padding, and number of channels. For fully connected layers, the crucial parameter is the number of units, also called the width of the layer.
Beyond the architecture used and data (pre-)processing, details about training parameters, loss function, optimizer, and regularization techniques are essential to replicate the results of such studies. While training parameters may not have as significant an impact as architecture parameters, they remain essential and can heavily impact training time and performance. Out of the 69 classification studies, 17 do not mention training parameters. The remaining studies generally mention learning rate, batch size, loss function, optimization method, and regularization techniques. Batch sizes, regularization techniques (dropout, L1 and L2 norm weight usage), number of epochs, early stop criteria, learning rates, and momentum parameters are crucial to watch out for. However, providing a detailed description for each parameter across all surveyed articles would not be very useful, as each study has its own setting, and no general rule can be extracted. However, further details on these parameters can be found in the complete table (https://tinyurl.com/ub3s5mr).
Among the studies that mentioned the optimization algorithm used for training their network, 29 studies used Adam [137]) or some variation thereof, while four used Stochastic Gradient Descent (SGD). Adam, short for Adaptive Moment Estimation, uses both the first-order and second-order moments of the gradient to adapt the learning rate, while SGD only uses the first-order moment. Adam typically converges faster than SGD.
As far as the loss function is concerned, most of the surveyed studies used cross-entropy to measure the discrepancy between the predicted probabilities and the actual distribution. More specifically, it computes the loss by taking the negative log of the probability assigned to the true class, aiming to minimize this loss to improve model accuracy.
Validation techniques are essential to training any model in a classification setting, enabling the proper evaluation of the model’s performance. In a setting where generalizing to new subjects is essential, such as BCI or CAD, it is capital to exclude multiple subjects’ data from the training process. This ensures the test of the trained algorithm’s generalization capabilities on entirely new and unseen data. It is also essential to evaluate and fine-tune the algorithm’s performances on training and validation datasets that do not share information with each other or the test set. Among the 69 studies in this category, 26 used the K-fold validation technique (K varying from two to ten), 28 used a simple train/valid/test set split, 10 used Leave One Out or Leave P out at the subject level. Five did not mention any validation technique in their analysis pipeline. The train/valid/test random splitting method is generally preferred in deep learning settings where the amount of data and the network size make training times very long, sometimes spanning multiple days. A large amount of data will dilute the split’s impact on the algorithm’s final performance. Note that when data is scarce, it is better to use bootstrapping or cross-validation to validate the model’s performance before evaluation on a test set, as it will lead to a less noisy estimate of performances ([138]). The Leave One Subject Out (LOSO) cross-validation strategy is preferred when subject generalization is essential, and the amount of available data is small. Note that when using K-fold or random split, it is crucial to think about whether it is important or not that the algorithm has seen parts of one subject’s data in the training set before validating or testing. Otherwise, the algorithm might end up identifying the subjects instead of solving the task it was designed for.
ANNs are powerful and can tackle a large set of problems; however, they present challenges in terms of interpretation, implementation, and tuning. In some instances, it is unnecessary to use ANNs for problems that simpler algorithms can solve. This is why it is essential when using ANNs to check whether we are using the right tool for the right problem by computing a baseline of performances that can be reached with more classical machine-learning approaches. Unlike ANNs, proper feature selection is essential for classical algorithms, particularly when handling high-dimensional like MEG data. Among the papers in the Classification category, 26 out of the 69 studies did not compare their ANN’s performance to simpler algorithms’ performances.
A critical part of decoding studies is understanding what aspect of the data allowed decoding. It is essential to find where and how information is encoded in the data to understand more about the brain. When using ML algorithms, especially ANNs, it can be hard to make sense of what is happening in the latent space. This is why using visualization tools will help with the interpretability of the network. Of the 69 papers in the classification category, only 17 include some interpretation or visualization techniques applied to the network. Only one of these study is a BCI sub-categorized study ([57]), two are event detection studies ([78, 84]), four are clinical sub-categorized studies ([68, 61, 64, 72]), and the remaining 10 are decoding studies ([26, 27, 34, 42, 32, 18, 44, 40, 41, 35]).
3.3 Modeling studies
The idea of using AI techniques, particularly deep learning, to reverse engineer brain function has turned into a thriving topic in Neuro-AI. In the context of this review, we refer to ’Modeling’ studies that use ANNs to build models of some function of the brain. In contrast to the papers that use ANNs to classify MEG data, the studies in this category do not use MEG data as input for ANNs. Instead, they primarily involve comparisons between the ANN activations and MEG recordings obtained in response to presenting the same visual or auditory stimuli to both systems (i.e. the ANN and the human brain). Therefore, the difference between the MEG-ANN modeling work and the ANN-powered MEG classification work is fundamental and conceptual. The underlying rationale for conducting such comparisons is the hypothesis that higher similarities between the artificial and biological neural responses indicate greater functional similarities between the ANN model and the neural network mechanisms.
Although the ANN training parameters, regularization, and validation techniques remain essential in this context for model training, they do not have the same type of impact on the results as for the classification studies discussed in the previous section. In the following, we’ll overview the main trends observed across the modeling studies, starting with the principle sensory processes that have been explored.
Four out of the 16 studies in this category investigated the Auditory cortex ([98, 99, 100, 101]), while nine focus on the visual cortex ([87, 88, 89, 90, 91, 92, 93, 96, 97]). Additionally, two studies aimed to model the visual word recognition in the human brain ([94, 95]) and two set out to model speech recognition ([102, 103]). Out of the eleven studies investigating the visual cortex (including visual word recognition), eight use Representational Similarity Analysis (RSA) or a variant RSA-based approach, one uses a method based on ’neural predictivity’ ([92]), and the last two use ’profile responses’ ([94, 95]). Among the studies focusing on modeling the function of the auditory cortex (including speech recognition), two comparing representations of ANNs to those of the brain when presented with auditory stimuli also use a variant of RSA ([101, 103]). Three employ methods based on neural predictivity ([98, 99, 102]). The last study employs ’Dimensionality Analysis’ ([100]).
3.3.1 Multivariate Pattern Analysis (MVPA)
Nine out of the 16 studies surveyed in this category begin with a Multivariate Pattern Analysis (MVPA) analysis ([87, 88, 89, 90, 93, 94, 95, 96], [100]). MVPA is a statistical approach in neuroimaging that uses standard machine learning algorithms, such as LDA or SVM, to analyze and interpret patterns of brain activity across multiple voxels or sensors, thereby facilitating the identification of cognitive states from detailed and multifaceted neural data. But why do many studies in this category begin by using MVPA? The primary objective of conducting MVPA analysis before initiating the similarity analysis between MEG and ANN is to help identify patterns and features in the data most relevant for distinguishing between different mental states or stimuli. Applying MVPA first effectively reduces the dimensionality of the MEG data. It brings focus to the most informative features, which can then be used to compare stimulus representations across artificial and biological networks.
It is helpful to note here that out of the studies which used MVPA, three actually use MVPA results (i.e. decoding accuracy) to build similarity measures ([89, 90, 100]). Furthermore, some studies describe the entire process of classification followed by similarity analysis as MVPA.
3.3.2 Modeling pipeline
Most of the articles included in the modeling category essentially compare neuromagnetic activity and the ANN activations generated by the same task (e.g., visual stimulus categorization).
To achieve this, MEG brain activity is typically recorded from healthy subjects as they engage in simple categorization or identification tasks involving either visual or auditory stimuli. Subsequently, these same stimuli are fed to an ANN model previously trained on a similar task and data distribution. In practice, different ANNs, such as CNNs for image stimuli and RNNs for audio stimuli, perform the same activities as the human subjects in the MEG study. This is then followed by a comparison of stimulus representation between ANNs and human brains. While some studies opt to train their ANN models from scratch, others use pre-trained models to bypass the resource-intensive training phase.
More concretely, once the models are trained, the same stimuli used for collecting brain data are fed to the ANNs and the layers’ responses’ are extracted. These responses are then considered as the ANNs’ activity, which will then be compared to the MEG activity induced by the same set of stimuli. That said, in one study, the authors concatenated the responses from all the layers responses to create a single model-level response [89]. Once the brain data are collected and the ANNs’ activations have been extracted, the next step is to compare them. In the following, we will review the main two methods used across the studies we surveyed: Representational Similarity Analysis (RSA) and neural predictivity. The other two remaining methods used in this context are ’Profile responses’ [94, 95] and ’Dimensionality analysis’ [100].
Representational Similarity Analysis (RSA): RSA is an established and widely used computational method to compare patterns of neural activity across different conditions or stimuli. This is achieved by constructing and analyzing similarity matrices representing the correlation or distance between the neural responses to each pair of conditions. These similarity matrices are known as Representational Dissimilarity Matrices (RDMs), which serve as a crucial tool in RSA by providing a quantitative measure of how neural responses differ across various experimental conditions, thus facilitating a deeper understanding of the underlying representational structures in the brain. In some studies, Representational Similarity Matrices (RSMs) are used instead of RDMs, employing direct correlation as the similarity metric rather than (1-correlation) for RDMs. Occasionally, decoding accuracies, an indirect measure of dissimilarity, are used instead of correlation measures ([89, 90, 100]).
The methodology for computing RDMs from fMRI data is generally uniform across many studies, typically focusing on Regions of Interest (ROI) RDMs. In this approach, the neural activity within a specified ROI is represented by a 1D vector describing the activity of each voxel. However, there are alternative methods that do not rely on predefined ROIs. One such alternative is the searchlight RSA (sRSA) technique, which involves specifying the shape and size of searchlight regions to scan the brain. Although computationally intensive, sRSA offers the advantage of making fewer assumptions about specific brain regions.
By contrast to RSA applications to fMRI data, the high temporal resolution of MEG provides the opportunity to conduct time-resolved RSA either in the sensor or in source space. To handle this additional time dimension, six of the 11 studies using RSA consider a 1D response vector for each time point by aggregating the activity of all brain sensors or specific ROIs’ voxels’ responses (in the case of source space). Ultimately, they generate time-resolved RDMs, sometimes referred to as ’RDM movies’ (see [91]). This particular RSA variant was termed RDA by [91]. Among the remaining studies employing RSA, two used a temporal variant from searchlight RSA, termed spatiotemporal sRSA (ssRSA) [101]. Instead of computing RDMs for every time point, the authors of these studies use time segments of a fixed length (25 ms for [101] and 16 ms for [87]). These studies combined this approach with searchlight RSA to construct multiple RDMs for different regions using a specific time interval. The final studies represented a combination of both approaches [93, 97, 103].
The eight studies computed the RDMs for the ANNs, one RDM per layer, with an additional RDM per model in one study.
Finally, a similarity score is computed by correlating the RDMs of MEG data with those of ANNs. The objective is to evaluate the similarity between the geometrical representations of MEG and ANNs across both time and space. For more information on RSA, see [139, 140, 141].
Neural predictivity: As an alternative approach to RSA, neural predictivity essentially involves using an ANNs’ layer activity to predict brain activity. This prediction is typically achieved using an ML algorithm, such as linear regression or a support vector machine. Some subjects are used to train the ML model, while the remaining subjects are kept aside as a test set. The algorithms’ performances are evaluated by comparing the predicted brain activity with the actual recorded brain activity, using correlation coefficients or mean squared error. We refer the readers to [142] for further details.
3.3.3 Estimating performance baselines
Assessing baseline performance is a crucial step in all analysis pipelines that seek to detect meaningful similarities between ANNs activity and MEG data. Generally speaking, determining baseline performance levels helps to distinguish the meaningfulness of the similarities observed between the ANN representations and the brain data, setting a benchmark for significance. Essentially, by establishing what can be expected by chance, or alternatively, what is maximally achievable given the noise in the data, we can confidently assert whether the similarities measured by RSA reflect genuine and meaningful correspondences between artificial and biological brain activity. In the following, we review the main methods of baseline performance used in the reviewed papers, focusing on Noise Ceilings (NC) and assessment of the untrained model’s performance.
Noise Ceiling (NC), or the shared-response model (SRM) comparison, refers to the theoretical upper limit or maximum level of similarity that can be achieved in the absence of measurement noise or variability in brain data. In other words, a noise ceiling provides an estimate of the best performance any model can achieve given the noise in the data. As such, it serves as a benchmark for the observed similarities between ANNs and the brain, indicating the extent to which individual subjects’ brain responses can be explained with a model-free approach. This concept often acts as a proxy for signal-to-noise ratio analysis, aiding researchers in interpreting the significance and reliability of their observed similarities by considering the inherent noise or variability in the data. As is common in RSA, the upper noise ceiling is estimated as the mean correlation between the group-average RSM and each participant-specific RSM. The lower noise ceiling is estimated as the mean correlation between the group-average RSM and each participant-specific RSM while iteratively excluding a given participant from the group-average. Some studies report both upper and lower noise ceilings. Although it is an essential measure, only five of the surveyed studies reported it. Among them, four used RSA as a similarity method ([87, 88, 91, 97]), one used neural predictivity ([99]) and the last used profile responses ([94, 95]). For detailed instructions on how to compute the noise ceiling, refer to [88].
Untrained models performance as baseline In RSA studies, a common approach to assess performance is to compare the results obtained with the trained ANN model to those obtained with an untrained ANN model. Some RSA-based studies (e.g. [143]) have found that even randomly initialized models can exhibit some similarity to the neural representation. Such an intrinsic similarity could result from built-in properties of specific models’ architecture (e.g., convolutional layers resembling the visual cortex). However, most RSA studies make implicit assumptions (or at least have some expectations) regarding the fact that trained ANNs would lead to enhanced ANN-brain similarities. In this context, comparisons of RSA results obtained with trained and untrained models can be very informative. Interestingly, out of the sixteen studies we found in this category, only three contrast their modeling results with those of a random model ([89, 94, 95, 99]).
While most studies focus on a single architecture, some include multiple ANN models. By conducting RSA analyses between the brain data and each one of the models, such studies can pick up the specific network architecture and training properties that increase the similarities between the ANN and biological responses. In [87], the authors employed a feature-based model where RDMs are computed based on the features extracted from the stimuli. In [89], the authors added a model trained on noise and an unecological model where images have been assigned random labels. Furthermore, in [88], the authors compared their similarities with those obtained by a semantic model. In [97], the authors use semantic models that measure semantic similarity between objects, i.e., the degree of resemblance in meaning between two pieces of text, such as words or sentences. In addition, it is common to train the same ANN on several similar tasks to assess the importance of the objective functions.
3.3.4 The data
Before we discuss the data used in the modeling papers, it is essential to recall that, unlike studies that use ANNs to classify MEG data, the research reviewed in this category does not input MEG data into ANNs. Instead, it primarily compares ANN activations with MEG recordings, observing responses to the same visual or auditory stimuli presented separately to both systems. For instance, in the study by Kietzmann et al. [91], the input data consists of images from different categories (animate or inanimate objects, faces, etc). Although MEG data are not used as inputs to the ANNs in these studies, they are still used to compare the representations in brain signals and the latent space variables of ANNs.
All MEG data collected for the studies in this category were preprocessed using a similar procedure, which in principle consists of a band-pass filter, an artifact removal and/or correction technique, and data down-sampling. The band-pass filtering’s lower and upper cutoff frequencies are 0.03 or 0.1Hz to 300 or 330Hz. However, there seems to be no discernible pattern explaining the sampling frequencies, the de-noising, or artifact removal techniques used (see 5). Among the 10 studies using RSA, five have used source reconstruction before computing RDMs ([87, 91, 101, 94, 95, 97]). It is worth noting that only two studies explored the frequency domain of MEG data by investigating the similarities in one or more frequency bands ([100, 99]).
3.3.5 The networks
The choice of ANN architecture is a particularly relevant aspect of the ANN-MEG modeling studies because the ultimate goal here is to assess and interpret similarities and discrepancies between the way information is processed in biological and artificial networks. As a general rule, the articles that focus on the visual cortex use a CNN architecture trained on image classification. Out of these studies, one incorporates lateral connections between layers, in addition to the typical top-down connections, introducing temporal relationships in their ANN ([91]). In this study, the ANN is trained to learn how to recreate MEG data RDMs from the stimuli image displayed to the subject. This is achieved through changing the ANN’s objective function. About half of the studies interested in the auditory cortex use LSTM ([100, 98]), the remaining studies use Transformer network and a CNN ([99]), an MLP architecture ([101]) or the BERT ([144]) model [103]. Finally, in [94, 95], the authors investigated the neural responses for a visual word recognition task by comparing the visual cortex’s response patterns to those of CNN architectures.
We found that while a few studies in this category train the networks from scratch, the majority (nine out of the 16) use pre-trained architectures from well-established types of networks, including VGG variants such as VGG-S (streamlined), VGG-F (fast), and VGG-11 (11 layers), which excel in image recognition. Other popular architectures include AlexNet, known for its effectiveness in image classification; CORnet-S, a brain-inspired model for predicting neural responses ([145]); and BERT, a transformer-based model widely used in natural language processing.
3.4 Other studies
This third category includes all the studies we found that combine MEG with ANNs but do not technically fall into the classification or modeling categories. Generally speaking, these studies cover a variety of topics such as data preprocessing techniques ([104, 105, 106, 107, 108, 109, 110, 111]), source estimation (solving the MEG inverse problem) ([128, 129, 130, 131, 132, 134, 135, 133, 146]) and pure methods-oriented articles ([113, 114, 115, 116, 117, 147, 124, 119, 118, 120, 127, 121, 123, 126, 125]).
Studies that employ ANN to tackle the MEG inverse problem (i.e. estimating the cortical sources from the sensor level data) typically employ simulated data from known sources, projecting them onto M/EEG sensor space with varying signal-to-noise ratios (SNRs). This simulated data is then used as input for an artificial neural network (ANN) pretrained to predict the original source location. Performance evaluation commonly involves assessing peak or average location error, often in conjunction with Area Under the Curve (AUC) metrics or spatial dispersion measures with AUC. Source estimation papers frequently use simulated data first to train their network, and then further validate their models on actual MEG data and compare their performance to other established techniques such as Minimum Norm Estimate (MNE) ([148]), Beamformer ([149]) or e- or s-LORETA ([150]). Studies in this subcategory exclusively use RNNs, ResNets, or CNNs, sometimes augmented with attention. Overall, the application of ANNs to source localization shows promise. Recently introduced novel baseline methods like ConvDip ([151]) for EEG data show that ANNs have the potential to yield innovative solutions to the complex inverse problem for MEG data.
Methodology-focused papers typically present innovative architectures or visualization tools optimized for MEG analysis. For instance, in ([125]), the authors propose a novel approach utilizing contrastive learning to develop a foundational model for aligning neural data and visual stimuli across multimodal brain activity representations. Additionally, in ([121]). The authors introduce two new classes of foundational models trained on MEG forecasting tasks, achieving state-of-the-art performance on decoding tasks. Note that methodology-focused studies may overlap with decoding-subcategorized studies but are distinguished by their broader focus on introducing innovative techniques to enhance decoding task efficiency rather than addressing a specific decoding problem.
Preprocessing-focused papers present innovative methods to augment traditional MEG analysis pipelines, targeting improvements in signal-to-noise ratio (SNR), artifact removal, and overall data quality. This subcategory includes a pioneering 2016 preprocessing study proposing Time Contrastive Learning (TCL), an unsupervised representation learning technique, as an alternative to classical Independent Component Analysis (ICA) for artifact removal ([109]). Another study [110] presents a highly accurate artifact classifier, leveraging a hybrid architecture combining deep 1-D and 2-D Convolutional Neural Networks (CNNs) to automate the detection and removal of pervasive whole-brain artifacts, including eye blinks, saccades, and cardiac artifacts. Furthermore, the authors in ([111]) propose a Denoising Autoencoder (DAE) to enhance MEG SNR, providing valuable neurophysiological insights while also offering a user-friendly tool for MEG data visualization and preprocessing.
Unfortunately, only nine studies out of the 34 papers surveyed in this category have their code publicly available or available upon request.
3.4.1 The data
Across the papers we reviewed in this category, the number of trials varied from 294 to 620256, and the number of subjects ranged from three to 676. Nine studies did not mention the amount of data used to train their networks, and two did not mention the number of participants. The studies mostly use MEG data as inputs, occasionally incorporating independent components (IC) of the signal, and two studies use connectivity graphs. Most studies focused on solving the source localization problem use anatomical MRI data in combination with the MEG data.
When MEG data is used, it is generally down-sampled to around 250Hz with some exceptions keeping the sampling rate high (2034Hz in [113], and 810Hz in [132]).
Out of the 34 studies classified as "Other", four used data augmentation ([113, 110, 107, 119]) and seven studies involved simulated data ([131, 130, 129, 134, 111, 133, 41]).
Interestingly, papers classified as methods studies only used MEG data to evaluate the performance of specific methods. It is key in such studies to evaluate the generalization potential of the proposed methods by testing them on different kinds of MEG data recordings (task, rest, etc) or MEG data from different datasets. Yet, we found that only nine out of 16 studies tested their method on more than one dataset or data type. ([117, 113, 115, 119, 121, 124, 123, 126, 125]).
3.4.2 The networks
Most Method and preprocessing studies use CNN-based architectures, while most Source localization-oriented studies use recurrent architectures (mostly LSTM). CNN-based architectures include simple CNNs ([105, 110, 107, 104, 108, 106, 116, 117, 114, 129, 120, 123, 124, 146, 125, 126]) but also Attention-Augmented CNNs ([113, 135]) and recurrent CNNs ([115]). The other studies use MLP ([105, 127, 129]), auto-encoders ([111]), CycleGAN ([119], ResNet ([133], or novel architectures such as Multiple Penalized State Space (MPSS) models ([134]).
Validation techniques are critical when it comes to evaluating the performance of novel methods. Depending on the dataset size, cross-validation might be a mandatory step to accurately assess a model’s prediction and generalization capabilities. All but one of the studies in this category mentioned some kind of validation technique. Source localization studies generally use a simple train/valid split that is sometimes bootstrapped. Method and preprocessing studies use a wider variety of validation techniques but can be condensed into either K-Fold or LOSO and simple train/valid/test splits.
Furthermore, we found that thirteen studies compared the performance of their methods to those of other simpler or more common alternatives ( [104, 105, 132, 129, 127, 131, 130, 122, 121, 123, 124, 135, 146]). Source localization studies generally compare the performance of their method to how Minimum Norm Estimate (MNE) or Beamformer would perform on the same data. Recent source localization studies also compare their results to e- or s-LORETA and ConvDip [151].
In this category, interpreting or visualizing the latent space of the networks might be less crucial for tasks like source localization or preprocessing. However, it remains vital in the methods subcategory, primarily when the study aims to introduce a new architectural approach for decoding tasks. As a matter of fact, in the method subcategory, nine out of 16 studies included some visualization of the latent space of their networks. All of them use a different technique: filter visualization ([113]), activation maps ([115]), channel and time contribution maps ([117]), feature visualization ([116]). In the preprocessing subcategory, five out of nine studies visualize at least partly the latent space of their network mainly through the use of grad-CAM ([108, 105, 110]). Interestingly, none of the source localization studies use any interpretation tool.
4 Discussion
4.1 Added value of combining ANN with MEG
ANNs are increasingly used alongside neuroimaging modalities such as fMRI, EEG, and MEG to enhance our understanding of brain function. fMRI is highly valued for its spatial resolution and ability to map cognitive states through blood flow changes. However, its slower temporal resolution limits its capacity to capture rapid neural dynamics. In contrast, both MEG and EEG offer the temporal precision necessary for analyzing fast brain activities. However, MEG distinguishes itself with superior spatial resolution and reduced susceptibility to scalp and skull distortions. Preprocessing steps, such as artifact removal and feature extraction, are crucial for ensuring that the high-dimensional MEG signals are effectively utilized by ANNs. These steps often form the foundation for downstream tasks by improving signal clarity and optimizing data input for architectures like CNNs, RNNs, and Transformers. In addition to preprocessing, ANNs are widely applied in MEG for classification tasks, including decoding, BCIs, clinical diagnostics, and event detection. By leveraging MEG’s temporal precision, ANNs enable detailed analyses of neural synchronization patterns, facilitating insights into cognitive processes such as perception, attention, language processing, and decision-making. For example, in neural classification, MEG’s precise temporal information supports real-time detection of complex states; in BCI applications, it enhances user control through responsive spatiotemporal decoding; in clinical diagnostics, it enables more accurate seizure detection by addressing signal distortions; and in event detection, it allows subtle oscillatory changes to be captured with higher fidelity. Beyond these applications, MEG–ANN integration plays a critical role in modeling studies, where the goal is to explore how MEG signals align with representations learned by neural networks. These studies compare MEG data with internal ANN activations, providing insights into how cortical activity corresponds to hierarchical network processing. MEG’s high temporal and spatial resolution provides a unique advantage in tracking dynamic neural activity, making it particularly well-suited for investigating how the temporal dynamics of ANN activations align with neural processes in the human brain, especially in domains like vision and audition, where precise timing is critical.
4.2 Current trends and dominant applications of ANNs in MEG
Taken together, the reviewed body of literature reveals a notable surge in publications and highlights the diverse applications of ANNs in MEG data analysis. This reflects not only a growing appreciation for data-driven approaches in neuroscience but also an increased interest in the unique contributions of ANNs to the field. Based on the papers we reviewed, it seems that the excitement about the potential of ANNs for advancing MEG research is distinct from the recent surge in using standard ML techniques. Our review shows that the power and versatility of deep learning is opening up frameworks for exploring MEG data that go beyond advanced statistical analyses. While many of the papers reviewed do indeed leverage the power of ANNs for classification tasks, many studies benefit from other strengths of ANNs. In particular, the reviewed literature shows that modeling studies that compare representations across ANNs and human MEG data seem to be gaining momentum. Methodological developments for MEG data analytics are also increasingly drawing on the strength and versatility of ANNs. Methodological developments for MEG data analytics are increasingly leveraging the strength and versatility of ANNs. This includes the introduction of novel architectures, enhanced data preprocessing techniques such as artifact detection and correction, innovative source reconstruction approaches, and the development of foundational models. These methodological innovations not only broaden the scope of MEG applications but also accelerate the pace of MEG research. That said, this review also highlights that there are still several critical limitations, including issues with interpretability and data scarcity, which will need to be addressed for the field to harness the potential of ANNs in MEG analysis fully. Further advancements - possibly including more robust foundation models for MEG - are necessary to improve model transparency and effectively manage limited datasets to achieve more reliable and generalizable results. In the next section, we will explore in greater detail the bottlenecks and challenges that emerge from the corpus of studies included in this review.
4.3 Current limitations and challenges
The diversity and wide range of ways ANNs have been applied to MEG research implies that the current challenges and limitations are also very diverse. Nevertheless, we will now attempt to categorize the dominant issues we observed across the reviewed papers into a handful of topics. Whenever possible, we offer strategic guidance to navigate these challenges in the hope of enhancing the robustness, efficacy, and reproducibility of future research in this rapidly evolving field. The wide range of ways ANNs have been applied to MEG has revealed a panorama of challenges and limitations as diverse as the methodologies themselves. In an attempt to synthesize for the reader the main challenges we identified across the reviewed literature, we categorize the predominant issues and pitfalls into a concise set of topics. First, we will address the general data scarcity problem prevalent in MEG-based neuroimaging studies. Next, we will underscore the importance of model validation techniques and the need for an accurate performance evaluation. We will then discuss the importance of hyperparameters in ANN-based MEG studies. Lastly, we will discuss the issue of reproducibility in ANN-based MEG studies.
4.3.1 Dealing with data scarcity
It is recommended that data augmentation techniques be used when feasible to cope with scarce data. For event-based epochs, introducing different event timings for each trial window will increase the sample size. In the case of continuous data, it is possible to use overlapping segments to artificially increase dataset size. Depending on the type of study and data, various data augmentation techniques can be added to a study’s preprocessing pipeline. For more details on the available methods, we refer the reader to [152], in which the techniques used for EEG data augmentation can also be applied to MEG data.
4.3.2 Model validation techniques
In many neural network applications, the vast amount of data combined with the typical training duration for ANNs makes cross-validation an impractical method for model validation. Instead, most of the time, the dataset is randomly split into three subsets to validate the model: training, validation, and test (also called evaluation set). The training set generally represents 50 to 80 percent of the initial dataset, and the remaining data is split in half to create the validation and test sets. The training set is employed to train the network with a set of hyperparameters and tested on the validation set. The test set is only used at the end when the architecture and hyper-parameters are fixed and once the desired performance has been reached on the validation set. It is solely used to evaluate the final model performance. However, when data is scarce, as it is often the case for neuroscience studies, it is recommended to use re-sampling methods such as bootstrapping or cross-validation to make use of the data as much as possible [138]. Cross-validation is often preferred (see Varoquaux’s response to [138]). It will allow the train and validation sets to be changed several times to acquire a less noisy measure of the model’s performance. In the studies included in this survey, when cross-validation is used, K-fold was the preferred technique, with the value of K ranging from 4 to 50. When the goal is to test generalization to new subjects, the "Leave P subjects out" cross-validation or a stratified version of K-Fold is recommended.
4.3.3 Model performance evaluation and baseline comparisons
Some classification studies did not include baseline results against which one could have compared the reported ANN performances. This can be a relevant concern given that classical shallow ML approaches (e.g. logistic regression ([153]), Random forest [154]), SVM ([155]), etc.) are generally easier and faster to implement than ANNs. Generally, one should always evaluate simple solutions before spending time and effort on more complex ones, which might also be more challenging to interpret. Similarly, only a few articles among the modeling studies mentioned some kind of baseline performance comparison. Ideally, modeling studies should report Noise Ceiling (NC) and/or evaluate the performance of their similarity analysis with an untrained model. Together, these measures would reinforce robust similarity scores’ reliability and help identify and discard weaker results.
4.3.4 Hyperparameter settings
In recent years, we have witnessed the development of increasingly deep networks aimed, for a large part, at enhancing performance in image classification tasks. However, the studies reviewed here reveal a trend toward relatively shallower network architectures. Intriguingly, when depth is evaluated as a hyperparameter in these studies, it appears to have a lesser impact on performance compared to other parameters, such as layer width or filter size. When using ANNs, the training phase is a pivotal component, and an incorrect selection of the optimizer and parameters could result in a network that fails to learn. Unfortunately, 16 of the classification studies did not provide enough information about the hyperparameters used. However, most studies specify the optimizer, learning rates, batch size, and the number of epochs, including the early stop condition when applicable. It is essential to systematically detail a model’s hyperparameters and training regimen (in any type of ANN study) to ensure the ability to reproduce results or to use the same model similarly with different data. In this context, a recommended practice is to also provide the source code alongside all the parameters and hyperparameters used to generate any given result of a study. Lastly, it’s worth mentioning that many studies in our review employ batch sizes that are not powers of two. However, using batch sizes that are powers of two is recommended to optimize computation and reduce training times [156].
4.3.5 The issue of reproducibility
In the field of neuroscience, concerns about the reproducibility of results are prevalent. The substantial costs and privacy issues related to acquiring neuroimaging or behavioral data restrict the availability of open-access datasets, thereby challenging the reproducibility of research conducted with such data ([157, 158]). In addition, this challenge is compounded by numerous studies that did not include enough information about the network architecture or the training parameters or did not provide access to their code, from which this information could be extracted. This lack of transparency further impedes the reproducibility of the results, making it difficult for researchers to validate and build on previous findings. The application of ML to neuroscience data faces challenges, mainly due to the relatively small/moderate sample sizes in neuroscience datasets. Besides, most algorithms require large amounts of data to achieve robust performance. Moreover, small datasets will make generalization harder to reach since having fewer data will further increase the impact of inter-subject variability. Solutions exist to tackle problems caused by small datasets, such as cross-validation, bootstrapping, and data augmentation. Still, more data will always be a preferable approach to achieve better results ([1]). However, the trend of small private datasets in the world of neuroscience is slowly changing with increasing awareness around the importance of open science practices, and public datasets are becoming more common.
4.3.6 Interpretation and visualizations
In Brain-Computer Interface (BCI) studies, where the primary concern is the algorithm’s performance, the black-box nature of ANNs is often considered less problematic. However, in many other applications, including brain decoding studies, gaining insight into what is happening in the model’s latent space is key to understanding how the information was extracted from the signal, as well as why and how decoding is possible. In domains like image classification or object detection, there are various tools (Grad-CAM, tSNE ([159]), UMAP ([160]) that allow the user to have a representation of what the network ’sees’ and what properties of the data are important in its decision-making process. A large proportion (52 out of 69) of the classification studies, where interpretation would be the most interesting, did not use these tools to gain insights into the representations learned and used by the model.
4.3.7 Note of caution on inconsistent nomenclature
Consistent terminology is, of course, vital for effective communication across all scientific disciplines. However, it becomes particularly crucial in emerging fields where the community still defines methodologies and core concepts. Adopting a uniform nomenclature is essential in a rapidly evolving research domain, such as the intersection of ANN and neuroimaging. Our extensive literature review reveals that terminology within classification studies is fairly consistent. However, we noted that the nomenclature in other categories, particularly within modeling studies, tends to show less consistency. To give a few examples, some research papers include RSA as a component of the MVPA analysis, while others do not. Additionally, some studies use neural predictability but do not explicitly mention it in their papers. In the modeling studies reviewed, the analysis pipelines exhibit considerable variety and flexibility, leading to notable diversity across the studies. For researchers looking to adopt these methodologies, we recommend a thorough examination of multiple studies. This approach will provide a comprehensive understanding of the various pipelines and help identify the methods most aligned with specific research objectives. Additionally, it is crucial to strive for consistency in terminology with the existing literature to ensure clarity and help the community move towards established procedures and nomenclature. For those interested in RSA in particular, insightful comments and critiques of the method can be found in [161].
5 Conclusion
Machine learning, notably the use of artificial neural networks, has become a cornerstone in contemporary research, reshaping both data analysis and modeling across diverse scientific fields. Although ANNs are relatively novel in the realm of neuroimaging, the collection of studies in this review demonstrates a rapidly growing interest in employing ANNs for both the analysis and modeling of MEG data. This trend is part of a broader movement towards integrating advanced data-driven and computational methods to deepen our understanding of neural mechanisms. The large body of work we reviewed indicates that ANNs are increasingly recognized as powerful tools in neuromagnetic imaging due to their robust performance, versatility, and ability to handle complex datasets. A comparison between our figure 1 and Figure 10b in [17] suggests a trajectory similar to that observed in the field of EEG, where ANNs have already shown extensive utility. This parallel points to a likely expansion in the number of studies employing ANNs for MEG data, further diversifying the applications and methodologies within the field.
One of the most remarkable observations is the incredibly diverse ways in which ANNs have so far been applied to MEG. While shallow ML tools like SVM or random forest classifiers are mainly applied to MEG in the context of brain decoding studies, ANNs open up a wide range of applications. As shown here, these extend from addressing the source estimation problem to providing new tools for MEG data cleaning and artifact rejection. Furthermore, our survey indicates an increasing interest within the neuroscience community in employing ANNs to build and evaluate information processing models in the human brain, reflecting the expanding role of these networks in advancing neuroscientific research [162].
Despite the remarkable advances in AI tools and the new opportunities that Artificial Neural Networks (ANNs) present for MEG data analysis and modeling, it is critical to avoid the allure of AI hype and the rush to employ complex AI algorithms solely for their novelty. Many research questions may still be best addressed using standard approaches that are well-established within the field and which often come with greater result interpretability. Moreover, the extensive range of parameters and hyperparameters involved in ANNs, coupled with the expertise required for their proper implementation, can sometimes introduce errors or yield misleading outcomes.
Regarding future developments, it seems that many exciting avenues are unfolding. To name just two of these: First, we believe that advances in AI interpretability and visualization tools will be key to enhancing the utility of ANNs in MEG research. Second, the recent trend towards AI foundation models in many disciplines, including neuroscience, will certainly also play an increasingly helpful role in MEG, both for basic and clinical applications. The scientific community is showing a growing interest in developing foundation models for brain imaging modalities ([163, 164]) and time series data ([165]). We anticipate that the neuroscience community will follow this trend ([121, 84]). Crucially, foundation models provide a convenient framework to combine multimodal data. Time will tell; one of the next paradigm shifts in this field will come from combining imaging modalities like MEG, EEG, and fMRI into a single multi-purpose foundation model.
Looking ahead, the future of this emerging research appears ripe with promising developments. Two particularly exciting prospects stand out. First, advances in AI interpretability and visualization tools are poised to significantly enhance the application of ANNs in MEG research, making these complex models more accessible and their findings more actionable. Second, the burgeoning adoption of AI foundation models across various disciplines, including neuroscience, is set to play a transformative role in both basic and clinical MEG applications. The scientific community’s growing enthusiasm for developing foundational models for brain imaging modalities ([163, 164]) and time series data ([165]) suggests that neuroscience will not be far behind ([121, 84]). These foundation models offer a versatile framework for integrating multimodal data, paving the way for a potential paradigm shift that could merge imaging modalities like MEG, EEG, and fMRI into a unified, multipurpose model. Integrating such modalities could revolutionize our understanding of the brain and its disorders.